OpenStudy (anonymous):

consider this math question, what place in the world can Fahrenheit an Celsius be equal????

OpenStudy (akshay_budhkar):

-40

OpenStudy (akshay_budhkar):

so a very cold place

OpenStudy (anonymous):

where temperature ix - 40

OpenStudy (akshay_budhkar):

ix?

OpenStudy (akshay_budhkar):

lol

OpenStudy (anonymous):

Siberia on a warm day

OpenStudy (akshay_budhkar):

warm day????

OpenStudy (anonymous):

warm for Siberia I mean!

OpenStudy (dumbcow):

F = 9/5C + 32 solve C = 9/5C +32 -4/5C = 32 C = 32*(-5/4) C = -40

OpenStudy (anonymous):

sorry akshay - just a silly joke!

OpenStudy (akshay_budhkar):

lol

OpenStudy (anonymous):

dumb cow was correct

OpenStudy (anonymous):

your math makes sense

OpenStudy (akshay_budhkar):

we all were correct

OpenStudy (akshay_budhkar):

lol

OpenStudy (anonymous):

i would love to know his?her credentials

OpenStudy (akshay_budhkar):

He is a master in maths

OpenStudy (akshay_budhkar):

what are yours?

OpenStudy (anonymous):

it is the norm, that a question is not answered by an other question

OpenStudy (anonymous):

what are your credentials? akshay

OpenStudy (akshay_budhkar):

LOl i am just kidding just trying to joke i am completely asleep now

OpenStudy (akshay_budhkar):

i have done my grade 12

OpenStudy (akshay_budhkar):

you?

OpenStudy (anonymous):

consider the haulting problem

OpenStudy (akshay_budhkar):

dun change the topic!lol what do u do?

OpenStudy (anonymous):

i am an ap high school student, and by the way i dint not intentionally change the topic, i left for a moment to do an errand, then i thought of a new question

OpenStudy (akshay_budhkar):

lol

OpenStudy (akshay_budhkar):

which grade?

OpenStudy (akshay_budhkar):

ya ask it

OpenStudy (anonymous):

10th

OpenStudy (akshay_budhkar):

cool

OpenStudy (anonymous):

ok i am going to post a new question

OpenStudy (akshay_budhkar):

ok

OpenStudy (anonymous):

think you can answer?

OpenStudy (akshay_budhkar):

truthfully no for today i am almost asleep!

OpenStudy (akshay_budhkar):

i can answer it tomorrow

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but then who?

OpenStudy (akshay_budhkar):

who? who am i?

OpenStudy (anonymous):

i mean who will anwer me question?

OpenStudy (akshay_budhkar):

you want it urgent? then post a new question simple

OpenStudy (akshay_budhkar):

ppl will answer

OpenStudy (anonymous):

ok thanks, i am rather new to this

OpenStudy (akshay_budhkar):

ohh

OpenStudy (akshay_budhkar):

wait

OpenStudy (akshay_budhkar):

do you know medal system?

OpenStudy (akshay_budhkar):

??

OpenStudy (anonymous):

no

OpenStudy (akshay_budhkar):

see where i have answered or replied on the right hand side can you see a good answer option? Click on it

OpenStudy (akshay_budhkar):

likee this

OpenStudy (anonymous):

would you be kind enough to explain to me why good sir/ma'am?

OpenStudy (akshay_budhkar):

you should hit correct answers

OpenStudy (anonymous):

do you honestly believe you merit a medal for your response ?

OpenStudy (akshay_budhkar):

lol its your option u liked it or not i give it to most of the ppl as i like and appreciate all d replies.. anyways.. bye tc

OpenStudy (anonymous):

bye